OpenStudy (anonymous):

In 1980, median family income was about $16,000, an in 2000 it was about $42,000. a) Find the slope of the line passing through the points (190,16000) and (2000, 42000). b) Intercept the slope as a rate of change. c) If this trend continues, estimate the median family income in 2003. Please help me and show me how to do this.

OpenStudy (anonymous):

ok, just give me a minute

OpenStudy (anonymous):

\[(190,16000) and (2000,42000)\]

OpenStudy (anonymous):

this is your given slope, ok

OpenStudy (anonymous):

then

OpenStudy (anonymous):

\[P1(190,16000) P2(2000,42000) that is P=point \]

OpenStudy (anonymous):

the formula for the slope or any slope is

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (anonymous):

then you just plug in those nubers

OpenStudy (anonymous):

P1(190,16000) P2(2000,42000) x1 y1 x2 y2

OpenStudy (anonymous):

y2-y1=42,000-16000 y2 y1

OpenStudy (anonymous):

and x2-x1=2000-190 x2 x1

OpenStudy (anonymous):

you subtract exactly what is on the equation

OpenStudy (anonymous):

42,000-16000 is 41840 and 2000-190 is 1810

OpenStudy (anonymous):

m=41840/1810 is m=32.11

OpenStudy (anonymous):

this is your slope of the line

OpenStudy (anonymous):

m=32.11

OpenStudy (anonymous):

m=y2-y1/x2-x1 witch is m=42,000-16000/2000-190